• Two equations of simple harmonic motion are given as follows : x₁= 4×10⁻² cos2π(t+1/8); x₂=3×10⁻² cos2π(t+1/4) . Find the equation of the resultant oscillation?
1
Expert's answer
2021-06-15T17:00:57-0400

Equation of SHMs are

"x\u2081= 4\u00d710^{-2} cos2\u03c0(t+1\/8); x\u2082=3\u00d710^{-2} cos2\u03c0(t+1\/4)"


Equation of resultant oscillation,

"y = x\u2081+x_2= 4\u00d710^{-2} cos2\u03c0(t+1\/8)+3\u00d710^{-2} cos2\u03c0(t+1\/4)"

"= [4\times 10^{-2}cos(2\pi t + \frac{\pi}{4})]+[3 \times 10^{-2}cos(2\pi t + \frac{\pi}{2})]"


Let "A = 4\times10^{-2} ; B = 3\times10^{-2}" to make equation simple.


"y = [Acos(2\pi t + \frac{\pi}{4})]+[Bcos(2\pi t + \frac{\pi}{2})]"


Expanding both cos functions,

"y = A[cos(2\pi t )cos( \frac{\pi}{4}) - sin(2\pi t )sin( \frac{\pi}{4})]+B[cos(2\pi t)cos(\frac{\pi}{2})-sin(2\pi t)sin(\frac{\pi}{2})]"


since, "cos(\pi / 4) = sin(\pi / 4) = \frac{1}{\sqrt{2}}; sin(\pi /2) = 1;cos(\pi /2) = 0"


"y = \frac{A}{\sqrt{2}} [cos(2\pi t) - sin(2\pi t)] - Bsin(2\pi t)"


"y = \frac{A}{\sqrt{2}}cos(2\pi t) - ( \frac{A}{\sqrt{2}} + B) sin(2\pi t)"


Let "\frac{A}{\sqrt{2}} = Psin\phi; (\frac{A}{\sqrt{2}} + B) = Pcos\phi"


Then "y = Psin(\phi) cos(2\pi t) -Pcos(\phi )sin(2\pi t) = Psin(\phi - 2\pi t)"

"y = -Psin(2\pi t - \phi)"


where, "\phi = tan^{-1}\frac{\frac{A}{\sqrt{2}}}{\frac{A}{\sqrt{2}} + B} = tan^{-1}\frac{A}{A + \sqrt{2}B} = tan^{-1}\frac{4\times 10^{-2}}{4\times 10^{-2} + 3\times 10^{-2}\sqrt{2}} = 25.89^\circ"

and

"P = \sqrt{(\frac{A}{\sqrt{2}})^2 + (\frac{A}{\sqrt{2}} + B)^2} = 6.47 \times 10^{-2}"


Hence equation will be,

"y = -6.47 \times 10^{-2}sin(2\pi t - \phi)"